A company issues a ten-year bond at par with a coupon rate of 6.5% paid semi-annually. The YTM at the beginning of the third year of the bond (8 years to maturity) is 9% . What is the new price of the bond?

a. $860

b. $1,203

c, $1,031

d. $1,000
Answer #1

Price of a bond is mathematically represented as: 1 1-- (1 + i) PEC. - +- м. (1 + i) where P is price of bond, with periodi

M = $1000, n = 8 * 2 = 16 semi-annual periods, C = 6.5%/2 * $1000 = $32.50 (semi-annually); i = 9%/2 = 4.5% (semi-annually)

1- P = 32.5 * - (1+0.045)16 a 0.045 1000 (1 + 0.045) 16

P = $365.11 + $494.47

P = $859.57

P = $860 (Option a)